OpenStudy (anonymous):

can someone please help me! Math each geometric sequence with the correct sum. a1 = 3, a8 = 384, r = 2 a1 = 80, n= 7, r = - 1/2 a1 = 5, n= 12, r = 3 a1 = 625, n= 5, r = 3/5 Choose... 215/4 1,328,600 630/3 765 1441

OpenStudy (anonymous):

ok lets do this since \[\Large S _{n}=\frac{a _{1}*(1-r^n)}{1-r}\]

OpenStudy (anonymous):

so for case one a1 = 3, a8 = 384, r = 2 here n=8 \[\Large S _{8}=\frac{3(1-(2)^8)}{1-2}=765\] for second one a1 = 80, n= 7, r = - 1/2 \[\Large S _{7}=\frac{80(1-(-1/2)^7}{1+1/2}=\frac{215}{4}\] can you do the rest by yourself ?

OpenStudy (anonymous):

@brendas